OpenStudy (anonymous):

Differentiate 2x^2+4^46+76-2^4 with respect to the variable x.

OpenStudy (anonymous):

@Jesstho.-.

OpenStudy (anonymous):

ARE YOU READY??!!!

OpenStudy (anonymous):

YASSSSS

OpenStudy (anonymous):

Multiply the exponent of x by 2 giving \[x^2\] The answer is \[x^2\] \[x^2 \] evaluates to \[x^2\]

OpenStudy (anonymous):

Multiply \[x^2\] and 2

OpenStudy (anonymous):

Multiply \[x^2\] and 1

OpenStudy (anonymous):

The \[x^2\] just get copied along

OpenStudy (anonymous):

The answer is \[x^2\]

OpenStudy (anonymous):

2*x^2 evaluates to \[2*x^2\]

OpenStudy (ipwnbunnies):

<_<

OpenStudy (anonymous):

4^46 evaluates to \[4^46\] \[4.95176\times10^+ {27}\]

OpenStudy (anonymous):

It's suppose to be 10 plus ^ 27 @ThatboiJayc;

OpenStudy (anonymous):

kk

OpenStudy (ipwnbunnies):

>_> The derivative is 4x. Derivative of constants are 0, so most of this goes to 0. Derivative of 2x^2 is 4x, using the power rule. \[\frac{d}{dx} [2x^2] = 2*2x^1 = 4x\]

OpenStudy (anonymous):

Yeah I know that. I was just showing him the steps, dude.

OpenStudy (anonymous):

I'll skip some steps @ThatboiJayc; Apply the "power rule" to x^2 (the derivative of x^n=nx^n-1)

OpenStudy (anonymous):

Multiply 2 and 2x Multiply 1 and x The x just gets copied along. x The answer is 4x

OpenStudy (anonymous):

Differentiating 2x^2 with respect to x

OpenStudy (anonymous):

4.95176x 10+^27 does not contain any reference to the variable x. So the derivative of this term with respect to x is zero (0) (it is a constant with respect to x).

OpenStudy (anonymous):

Oh I see

OpenStudy (anonymous):

Good, If you need anymore help with an problem just tag me :)

OpenStudy (anonymous):

I understand quite well and thanks :p